OpenStudy (ladiesman217):

i give medals and fans. What are the characteristics of a radical equation? (1 point) How is solving radical equations similar to solving linear equations? (1 point) Why is it important to check the solutions to a radical equation? (1 point) Create a radical equation. Describe in complete sentences and demonstrate the process in finding its solution(s). (2 points)

OpenStudy (anonymous):

1) the main characteristic is that the variable is under the radical. 2) when you solve a linear equation, you isolate the variable ( try to get x alone on one side of the equation). When you solve a radical equation, you try to get the radical alone on one side. 3) usually, you solve a radical equation by raising it to a power. Since second degree, or higher degree, equations have more solutions than first degree, you may be increasing the number of solutions. These are called extraneous roots. You have to throw these out. 4) easy or hard? 5√(x-2) + 7= 22...... first subtract 7 5√(x-2) = 15..... next divide by 5 √(x-2)= 3..... now square both sides X-2= 9.....add 2 X=11 Check! 5√(11-2)+7= 22 5√9 +7= 22 5*3+7=22 I used: http://answers.yahoo.com/question/index?qid=20111229155211AATFNHV

OpenStudy (ladiesman217):

ok thnx